OpenStudy (anonymous):

A 71.0kg fireman slides down a 4.73 meter pole by applying 350. N against the pole with his hands. If he slides from rest, how fast is he moving when he reaches the ground?

OpenStudy (anonymous):

taking gravity as 9.8m/s^2 weight of man = 71.0*9.8 = 695.8N Resultant force = weight of man - force the man applies = 695.8 - 350 = 345.8N total amount of energy = Force*distance = 345.8*4.73 = 1636J KE = 1636J = 0.5*m*v^2 therefore, speed = (1636/30.5)^0.5 = 7.32 m/s

OpenStudy (anonymous):

i dont know if this is correct because i am only in middle school

OpenStudy (anonymous):

I think you are correct.

OpenStudy (anonymous):

oh thanks